OpenStudy (anonymous):

Penicillin has a half life of 36 hours in the bloodstream. Suppose that a patient recieves an initial dose of 30 milligrams of penicillin at 9am. How much penicillin will be in the patients blood at 9pm that night? Approximate when the penicillin concnetration will reach 1% of its original level? What is the decay rate for penicillin?

OpenStudy (turingtest):

I'm not always great at these kind of questions, but the rate change of the drug \(dP/dt\) is proportional to some constant \(r\) times the amount of the drug \(P\) because it's decay it's negative... so we can write\[\frac{dP}{dt}=-rP\]

OpenStudy (turingtest):

the solution to that differential equation is\[P(t)=P_0e^{-rt}\]where \(P_0\) is the initial dose. I think we should say \(t\) is in hours now we need to find \(r\)...

OpenStudy (turingtest):

we are given that the half life is 36 hours you say? not minutes, right?

OpenStudy (anonymous):

Correct

OpenStudy (turingtest):

ok then we have that the amount in the bloodstream is half when\[\frac{P_0}2=P_0e^{-36r}\]can you solve this for \(r\) ?

OpenStudy (turingtest):

and since the intial dose \(P_0\) is known to be 30mg we can write that in if you like...\[\large P(36)=15=30e^{-36r}\]can you solve it for \(r\) ?

OpenStudy (anonymous):

6.662463305E-54

OpenStudy (turingtest):

that sounds wrong...

OpenStudy (turingtest):

\[\frac{P_0}2=P_0e^{-36r}\]\[\frac12=e^{-36r}\]\[\ln\frac12=-\ln2=-36r\]\[r=\frac{\ln2}{36}\]

OpenStudy (turingtest):

so from 9am to 9pm how many hours is that?

OpenStudy (anonymous):

36-12=24hours

OpenStudy (turingtest):

12 hours

OpenStudy (turingtest):

But aside from that, I've made a mistake... these really aren't my kind of problems, I was just trying to give you quick help now I need to figure out what I did wrong

OpenStudy (anonymous):

Thanks anyway...these aren't easy and I am finding difficulty imputing these formulas in my caluculator.

OpenStudy (turingtest):

ah, you want to do it the calculator way, eh? I see my mistake now though, the formula is...

OpenStudy (anonymous):

Our teacher allows calculator on the quiz...

OpenStudy (turingtest):

\[\huge P(t)=(30)2^{-\frac t{36}}\]a) plug in t=12 hours and you will have the amount in the body at that time b) solve\[\huge\frac{30}{100}=(30)2^{-\frac t{36}}\]for t and you will find when the concentration is one percent its original good luck!

OpenStudy (turingtest):

oh, and c) the decay rate is \(r\), which we found earlier

OpenStudy (anonymous):

(30)2^-13/36?

OpenStudy (turingtest):

from 9am to 9pm is 12 hours, so you need to put in t=12, not 13

OpenStudy (anonymous):

(30)2^-12/36=syntax error